# A POD-TANN approach for the multiscale modeling of materials and macroelement derivation in geomechanics

### v1

@@ 2 Theoretical background
At the core of the Coleman and Gurtin’s approach is the use of the second law, in the form of the Clausius-Duhem inequality. This is exploited, together with the first law, to obtain restrictions on the form of the constitutive equations. A general review of the theory may be found in [7]. For the sake of simplicity, isothermal processes and infinitesimal strain regime will be considered in this work. The relevant quantities are introduced as follows: $\boldsymbol{\epsilon}=Sym\left(\nabla\boldsymbol{u}\right)$, $p^{in}=\boldsymbol{\sigma}:\dot{\boldsymbol{\epsilon}}$, with $\boldsymbol{u}$ being the displacement field, $\boldsymbol{\sigma}$ the Cauchy stress tensor and $\boldsymbol{\epsilon}$ the infinitesimal strain tensor.
@@
$\boldsymbol{\sigma}-\frac{\partial\hat{\psi}}{\partial\boldsymbol{\epsilon}}=0,\quad d=-\frac{\partial\hat{\psi}}{\partial\boldsymbol{z}}\cdot\dot{\boldsymbol{z}}\geq 0.$ (1)
Furthermore, depending on the setting chosen by the modeler, Fenchel-Legendre transformations of the potential energy density are introduced so to switch from dependent and independent static and kinematic variables. To aid the formulation used in the following secitons, the transform between the Helmholtz and Gibbs free energy densities is recalled:
@@ 2.1 Formulation for multiscale homogenization problems
The formal derivation of the constitutive restrictions obtained in [22] is retained when a macroscopic constitutive equation is derived by homogenization of a microscopic system under prescribed boundary conditions. This has been demonstrated in [8], considering the framework of the asymptotic (espansion) homogenisation (AEH) theory, in incremental formulation, see, e.g., [24], [25], [26].
@@
in which the following volume average quantities are defined: $\boldsymbol{\Sigma}(\boldsymbol{x})\triangleq\left\langle\boldsymbol{\sigma}^{e}(\boldsymbol{x},\boldsymbol{y})\right\rangle$, $\boldsymbol{E}(\boldsymbol{x},\boldsymbol{y})\triangleq\left\langle\boldsymbol{\epsilon}^{e}(\boldsymbol{x},\boldsymbol{y})\right\rangle$, $\Psi(\boldsymbol{x},\boldsymbol{y})\triangleq\left\langle\psi^{e}(\boldsymbol{x},\boldsymbol{y})\right\rangle$, $D(\boldsymbol{x},\boldsymbol{y})\triangleq\left\langle d^{e}(\boldsymbol{x},\boldsymbol{y})\right\rangle$.
In the above equation, the macroscopic vector of internal state variables has been expressed using the symbol $\boldsymbol{\mathcal{Z}}$.
@@ 3 Thermodynamic-based Artificial Neural Networks for the multiscale problem
Relying on the theoretical background presented in Section 2, Masi et al. [7] have developed the so called Thermodynamics-based Artificial Neural Networks - TANN, to learn the constitutive behaviour of inelastic homogeneous materials. The initial TANN formulation has been further extended from homogeneity, to discrete heterogeneous micro-structured inelastic materials in [8].
@@
The weights are used to assess the relative output influence on the training process, the training losses, $\ell^{i}$ with $i=[\psi,\boldsymbol{\sigma},d,\{d\}]$, are computed using a metric of the error between the output and the training data.
@@ 3.1 Discovery of internal state variables using model reduction
Mathematically, the POD allows to write a generic field $\boldsymbol{u}(\boldsymbol{x},t)$ by means of an eigenfunction expansion (see, e.g., Lumley [10]):
in which the column of $\boldsymbol{\Phi}$, $\boldsymbol{\Phi}_{k}(\boldsymbol{x})$, are deterministic time-independent spatial functions (POD modes), modulated by the time-dependent coefficients $c_{k}(t)$. In practice, the optimal basis of modes is found exploiting the Singular Value Decomposition (SVD). Denoting with $\boldsymbol{A}$ a generic snapshot matrix of $\boldsymbol{a}$, the low-rank SVD truncation of $\boldsymbol{A}$, $\tilde{\boldsymbol{A}}=\tilde{\boldsymbol{U}}\tilde{\boldsymbol{S}}\tilde{\boldsymbol{V}}^{*}$, is used to identify the time-varying reduced order state variables of the macro-system:
$\tilde{\boldsymbol{A}}=\tilde{\boldsymbol{U}}\tilde{\boldsymbol{S}}\tilde{\boldsymbol{V}}^{*}=\boldsymbol{\Phi}\boldsymbol{C}\to\begin{cases}\boldsymbol{\Phi}=\tilde{\boldsymbol{U}},\\ \boldsymbol{C}=\tilde{\boldsymbol{S}}\tilde{\boldsymbol{V}}^{*}=\tilde{\boldsymbol{U}}^{*}\tilde{\boldsymbol{A}}.\end{cases}$ (9)
The matrix $\boldsymbol{C}$ collects the time-varying POD coefficients describing the influence of each POD mode at the considered instant of time.
@@ 4 Applications
This section will examine three applications of the proposed approach. In sections 4.1 and 4.2, consideration will be given to learning the homogenized behaviour of continuum microstructured periodic unit cells featuring inelastic behaviour. Then, in section 4.3, the data-driven derivation of a macroelement for the horizontal response of a monopile in a clayey layer will be detailed.
@@ 4.1 RUC with ellipsoidal inclusion
The first example focuses on a RUC with an ellipsoidal inclusion. The matrix and the inclusion obey an elasto-plastic model with isotropic hardening and Drucker-Prager yield criterion. Material parameters are reported in Table 1.
The training dataset has been generated by imposing random strain increments to the RUC, starting from a compression volumetric strain of -5e-4. This is done in order to prevent obtaining data in the positive (tensile) volumetric domain, which is seldom relevant for applications in geomechanics. The strain paths are created by sampling the increments from a standard normal distribution with a mean of 0.0 and standard deviation of 5e-4. In the sampling procedure, a limit for the second deviatoric strain invariant has been set to 0.015, sufficient to have more than half of the stress points in the plastic regime.
Strain paths of 1000 increments have been considered. The initial dataset results by the composition of five random strain paths. Then, the final dataset has been obtained by augmentation of the initial one. The procedure consists in rotating second order tensors of the microscopic stress and strain fields with a randomly sampled 3D rotation matrix, $\boldsymbol{R}$, while keeping unchanged invariant quantities, such as the energies. In addition to save computational time for dataset generation, this procedure allows the neural network to learn objectivity from data.
@@
ICs results by the collection of elastic and plastic microscopic strains and the maximum second invariant of the plastic strains at the Gauss points of the computational model at all increments, namely $\boldsymbol{\xi}=\left[\boldsymbol{\epsilon}^{el}_{i},\boldsymbol{\epsilon}^{pl}_{i},J_{2,max}^{pl}\right]$, at all Gauss points. Since the hardening is isotropic and linear in this example and there is no softening, $J_{2,max}^{pl}$ is the hardening variable, see [33]. These microscopic state variables are sufficient to describe the microscopic state at each Gauss point of the RUC.
@@
The IC reduction of dimensionality prior to the training of the energy network has been quantified in terms of a compression ratio, $CR$, defined as follows:
@@
Considering 25 POD modes, the compression ratio results in $CR=99.98\%$. This significant reduction is achievable due to the redundancy of information stored within the set of internal configurations (ICs). While the heterogeneity of the material induces variability in the strain fields, gradients predominantly occur in the region of the inclusion. Numerous Gauss points in the matrix, far from the inclusion, exhibit strain field values that change only slightly relative to each other, allowing for high compression.
To train the Helmholtz free energy network a supervised learning procedure has been adopted, collecting inputs and outputs as detailed in Section 3. All the inputs and outputs have been normalized to range between -1 and 1, with the exception of the Helmholtz energy and the rate of mechanical dissipation, for which positive and normalized values between 0 and 1 have been used.
The energy network is a feed forward net with a single hidden layer of 100 neurons with quadratic activation function of the form $p^{(k)}_{j}=\mathcal{A}^{k}(r^{(k)}_{j})=r^{(k)2}_{j}$, and a linear single output layer with biases set to zero. The latter guarantees the reference energy to be zero at an initial zero-valued state.
@@ 4.2 RUC with leaf-shaped inclusion
Figure 5-c presents the learning curves obtained during the training process. Remarkably, the neural network training converges after only 1,000 epochs, indicating a promising result. The figure displays the learning curves for 15, 20, and 25 POD modes. A significant improvement is observed from 15 to 20 modes, with the plateau value decreasing by almost one order of magnitude. This improvement is also evident in Figure 5-b, where the average energy reconstruction error and its standard deviation decrease by approximately one order of magnitude. Beyond 20 POD modes, the error curve’s slope remains nearly constant as the number of modes increases. This is reflected in the minimal difference between the stress loss curves for 20 and 25 modes, although a further decrease in the plateau value is noted with 25 modes. With 25 modes, the compression ratio is $CR=99.97\%$.
@@ 4.3 Macroelement of a monopile in saturated clay
Traditionally, the evaluation of the response of these types of foundations has been based on simplified approaches, such as those of p-y curves, as in [35] and [36], and more rarely on complex numerical analyses of the single pile and the foundation soil. Alternatives to these approaches are macro-elements. These macro-elements maintain the simplicity of use found in p-y curves, but they allow for the direct consideration of complex aspects of soil-structure interaction on the macroscopic response of the system, such as the progressive degradation of the soil’s stiffness and resistance, friction between the pile and the ground, and the hysteretic damping response of the soil, among others.
In this section, a macro-element capable of replicating the horizontal undrained cyclic response in terms of force-displacement of a monopile in a saturated clay layer is presented. The macro-element results from training a TANN on a numerical database obtained from finite element simulations of the macroscopic system.
@@
The TANN is combined with Proper Orthogonal Decomposition and applied to a macroscopic geotechnical system. Eventually, the studied 3D system is condensed into a macro-element with one degree of freedom.
@@ 4.3.2 Data generation and training of the TANN
The soil-pile system is reduced to a macro-element with one degree of freedom, considering the horizontal displacement of the point at the pile head positioned on the pile axis, $U_{H}$, as a function of the resultant of the horizontal forces at the pile head, $F_{H}$.
@@
As already shown in the preceding sections, the threshold of 1e-4 allows to achieve satisfactory results both in terms of accuracy in the reproduced response of the 1D equivalent system and of the reconstruction error of the microscopic fields.
Indeed, for this kind of applications the possibility of reconstructing the fields used as internal coordinates is particularly appealing. One of the major drawbacks of the classical macroelement formulations is the impossibility of retrieving microscopic information, once the macroscopic behaviour has been simulated. This poses clear difficulties in the engineering design, since often structural components are encapsulated by the macro-material whose response is of interest. Thus, an additional model is needed to link the global response of the classical macroelement to that of the structural components. As shown above, this in not needed with the proposed POD-TANN approach.
@@
Figure 13: TANN predictions in inference of a straining path outside the training set for cycles of positive and negative amplitude (a) and asymmetric cycles (b).

### v2

@@ 2 Theoretical background
At the core of the Coleman and Gurtin’s approach is the use of the second law, in the form of the Clausius-Duhem inequality. This is exploited, together with the first law, to obtain restrictions on the form of the constitutive equations. A general review of the theory may be found in [7]. For the sake of simplicity, isothermal processes and infinitesimal strain regime will be considered in this work. The relevant quantities are introduced as follows: $\boldsymbol{\epsilon}=Sym\left(\nabla\boldsymbol{u}\right)$, $p^{in}=\boldsymbol{\sigma}:\dot{\boldsymbol{\epsilon}}$, with $\boldsymbol{u}$ being the displacement field, $\boldsymbol{\sigma}$ the Cauchy stress tensor, $\boldsymbol{\epsilon}$ the infinitesimal strain tensor, $p^{in}$ the internal power, $\nabla$ the gradient operator and $Sym(\bullet)$ the symmetric part operator.
@@
$\begin{split}&(I\quad\text{Law})\quad p^{in}=\dot{\psi}+d;\\ &(II\quad\text{Law})\quad d\geq 0;\\ &(\text{Constraints})\quad\boldsymbol{\sigma}-\frac{\partial\hat{\psi}}{\partial\boldsymbol{\epsilon}}=0,\quad d=-\frac{\partial\hat{\psi}}{\partial\boldsymbol{z}}\cdot\dot{\boldsymbol{z}}\geq 0.\end{split}$ (1)
In the above expressions, $\psi=\hat{\psi}(\boldsymbol{\epsilon},\boldsymbol{z})$ is the Helmholtz free energy density, $d$ is the rate of mechanical dissipation. Furthermore, depending on the setting chosen by the modeler, Fenchel-Legendre transformations of the potential energy density are introduced so to switch from dependent and independent static and kinematic variables. To aid the formulation used in the following secitons, the transform between the Helmholtz and Gibbs free energy densities is recalled:
@@ 2.1 Formulation for multiscale homogenization problems
The formal derivation of the constitutive restrictions obtained in [22] is retained when a macroscopic constitutive equation is derived by homogenization of a microscopic system under prescribed boundary conditions. This has been demonstrated in [8], considering the framework of the asymptotic (espansion) homogenization (AEH) theory, in incremental formulation, see, e.g., [24], [25], [26].
@@
in which the following volume average quantities are defined: the macroscopic Cauchy stress tensor, $\boldsymbol{\Sigma}(\boldsymbol{x})\triangleq\left\langle\boldsymbol{\sigma}^{e}(\boldsymbol{x},\boldsymbol{y})\right\rangle$, the macroscopic small strain tensor $\boldsymbol{E}(\boldsymbol{x},\boldsymbol{y})\triangleq\left\langle\boldsymbol{\epsilon}^{e}(\boldsymbol{x},\boldsymbol{y})\right\rangle$, the macroscopic Helmholtz free energy density $\Psi(\boldsymbol{x},\boldsymbol{y})\triangleq\left\langle\psi^{e}(\boldsymbol{x},\boldsymbol{y})\right\rangle$, the macroscopic rate of mechanical dissipation $D(\boldsymbol{x},\boldsymbol{y})\triangleq\left\langle d^{e}(\boldsymbol{x},\boldsymbol{y})\right\rangle$, the macroscopic vector of internal state variables $\boldsymbol{\mathcal{Z}}$.
@@ 3 Thermodynamic-based Artificial Neural Networks for the multiscale problem
Relying on the theoretical background presented in Section 2, Masi et al. [7] have developed the so called Thermodynamics-based Artificial Neural Networks - TANN, to learn the constitutive behavior of inelastic homogeneous materials. The initial TANN formulation has been further extended from homogeneity, to discrete heterogeneous micro-structured inelastic materials in [8].
@@
The weights, $w^{i}$ with $i$ in $[\psi,\boldsymbol{\sigma},d,\{d\}]$, are used to assess the relative output influence on the training process; the training losses, $\ell^{i}$, are computed using a metric of the error between the output and the training data.
@@ 3.1 Discovery of internal state variables using model reduction
In the proposed POD-TANN framework, the POD is employed to reduce the dimensionality of the internal coordinates, $\boldsymbol{\xi}$, which are responsible for tracking the irreversible processes occurring at the microscale. By applying POD to the snapshot matrix $\boldsymbol{\Xi}$, constructed from successive snapshots of $\boldsymbol{\xi}$, the ICs are approximated using a set of optimal spatial modes, $\boldsymbol{\Phi}$, and time-varying coefficients, $\boldsymbol{\mathcal{Z}}$.
Mathematically, the field of internal coordinates, $\boldsymbol{\xi}$, can be expressed as:
$\boldsymbol{\xi}(\boldsymbol{x},t)=\sum_{k=1}^{r}z_{k}(t)\boldsymbol{\Phi}_{k}(\boldsymbol{x})=\boldsymbol{\Phi}(\boldsymbol{x})\boldsymbol{Z}(t),$ (8)
where $r$ is the number of retained POD modes, $\boldsymbol{\Phi}_{k}(\boldsymbol{x})$ are the time-independent spatial modes, and $\boldsymbol{Z}(t)$ is the collection of vectors of time-dependent POD coefficients, representing the reduced-order internal state variables (ISVs) at each time step.
The optimal POD basis, $\boldsymbol{\Phi}$, is derived through the Singular Value Decomposition (SVD) of the snapshot matrix $\boldsymbol{\Xi}$, which organizes the snapshots of the internal coordinate field. Through SVD, $\boldsymbol{\Xi}$ is approximated and decomposed as:
$\boldsymbol{\Xi}\approx\tilde{\boldsymbol{\Xi}}=\tilde{\boldsymbol{U}}\tilde{\boldsymbol{S}}\tilde{\boldsymbol{V}}^{*},$ (9)
where $\tilde{\boldsymbol{U}}$ contains the first $r$ time-invariant spatial modes (POD modes) of $\boldsymbol{U}$, $\tilde{\boldsymbol{S}}$ holds the retained singular values, and $\tilde{\boldsymbol{V}}^{*}$ contains the time coefficients, with * indicating the transposed-conjugate. By retaining only the dominant $r$ modes, the dimensionality of the ICs is reduced, yielding a compact yet representative set of ISVs, $\boldsymbol{\mathcal{Z}}$, that describe the system’s evolution. $\boldsymbol{\mathcal{Z}}$ is obtained projecting $\boldsymbol{\xi}$ onto the basis defined by the orthogonal retained modes, namely $\boldsymbol{\mathcal{Z}}=\tilde{\boldsymbol{U}}^{*}\boldsymbol{\xi}$ (see Appendix A for a detailed derivation). Exploiting the time-invariance of the POD-modes, the rate of change of the ISV are also easily found from the rate of change of the ICs, as follows:
$\dot{\boldsymbol{\mathcal{Z}}}=\frac{d}{dt}(\tilde{\boldsymbol{U}}^{*}\boldsymbol{\xi})=\tilde{\boldsymbol{U}}^{*}\dot{\boldsymbol{\xi}}.$ (10)
@@ 4 Applications
This section will examine three applications of the proposed approach. Sections 4.1 and 4.2 will focus on learning the homogenized behavior of continuum microstructured periodic unit cells exhibiting inelastic behavior. Then, Section 4.3 will detail the data-driven derivation of a macroelement for the horizontal response of a monopile in a clayey layer.
@@ 4.1 RUC with ellipsoidal inclusion
The first example focuses on an RUC with an ellipsoidal inclusion. Both the matrix and inclusion follow an elasto-plastic model with isotropic hardening and the Drucker-Prager yield criterion. The material parameters are provided in Table 1.
4.1.1 Data generation
The training dataset was generated by applying random strain increments to the RUC, starting from an initial compression volumetric strain of -5e-4. This approach was chosen to avoid generating data in the positive (tensile) volumetric domain, which is rarely relevant for geomechanics applications. The strain paths were created by sampling increments from a standard normal distribution with a mean of 0.0 and a standard deviation of 5e-4. To ensure that a significant portion of the stress points fall within the plastic regime, a limit of 0.015 was imposed on the second deviatoric strain invariant.
Strain paths consisting of 1000 increments were considered. The initial dataset was composed of five random strain paths. The final dataset was then obtained through data augmentation. This involved rotating the second-order tensors of the microscopic stress and strain fields using a randomly sampled 3D rotation matrix, $\boldsymbol{R}$, while preserving invariant quantities such as energy. This augmentation not only reduced computational time for dataset generation but also allowed the neural network to learn objectivity from the data.
@@ 4.1.1 Data generation
The ICs are obtained by collecting the elastic and plastic microscopic strains, along with the maximum second invariant of the plastic strains at the Gauss points of the computational model at each increment, expressed as $\boldsymbol{\xi}=\left[\boldsymbol{\epsilon}^{el}_{i},\boldsymbol{\epsilon}^{pl}_{i},J_{2,max}^{pl}\right]$ at all Gauss points. Since the hardening in this example is isotropic and linear, and there is no softening, $J_{2,max}^{pl}$ serves as the hardening variable (see [33]). These microscopic state variables are sufficient to fully describe the microscopic state at each Gauss point of the RUC.
@@
The reduction in dimensionality of the ICs prior to training the energy network has been quantified using a compression ratio, $CR$, defined as follows:
@@
With 25 POD modes, the compression is $CR=99.98\%$. This substantial reduction is achievable due to the redundancy in the information contained within the set of internal coordinates. Although the material’s heterogeneity introduces variability in the strain fields, the gradients are primarily concentrated around the inclusion. Many Gauss points within the matrix, especially those far from the inclusion, exhibit strain field values that change marginally relative to one another, enabling such high compression.
4.1.2 Training of the TANN
To train the Helmholtz free energy network a supervised learning procedure has been adopted, collecting inputs and outputs as detailed in Section 3. All the inputs and outputs have been normalized to range between -1 and 1, with the exception of the Helmholtz energy and the rate of mechanical dissipation, for which positive and normalized values between 0 and 1 have been used. As schematically depicted in Figure 1, the input of the network has variable dimensionality, depending on the number of retained POD modes, in agreement with the following expression: $6+2\cdot N_{POD}$. Despite the latter variability, the training of the TANN has been performed keeping constant the number of hidden layers and neurons of the energy network, as detailed below.
The energy network is a feed forward net with a single hidden layer of 100 neurons with quadratic activation function of the form $p^{(k)}_{j}=\mathcal{A}^{k}(r^{(k)}_{j})=r^{(k)2}_{j}$, and a linear single output layer with biases set to zero. The latter guarantees the reference energy to be zero at an initial zero-valued state. The Nadam optimizer has been adopted, see [34], with a learning rate of 5e-5. Weight and biases have been initialized using the Glorot uniform initializer, see [35]. A constant mini-batch size of 1000 has been used.
@@ 4.2.1 Data generation
4.2.2 Training of the TANN
The training of the TANN has been achieved using the same setting detailed in section 4.1.2. Figure 5-c presents the learning curves obtained during the training process. Remarkably, the neural network training converges after only 1000 epochs. The figure displays the learning curves for 15, 20, and 25 POD modes. A significant improvement is observed from 15 to 20 modes, with the plateau value decreasing by almost one order of magnitude. This improvement is also evident in Figure 5-b, where the average energy reconstruction error and its standard deviation decrease by approximately one order of magnitude. Beyond 20 POD modes, the error curve’s slope remains nearly constant as the number of modes increases. This is reflected in the minimal difference between the stress loss curves for 20 and 25 modes, although a further decrease in the plateau value is noted with 25 modes. With 25 modes, the compression ratio is $CR=99.97\%$.
4.2.3 Inference of the TANN
@@ 4.3 Macroelement of a monopile in saturated clay
Traditionally, the evaluation of the response of these types of foundations has been based on simplified approaches, such as those of p-y curves, as in [37] and [38], and more rarely on complex numerical analyses of the single pile and the foundation soil. Alternatives to these approaches are macroelements. These macroelements maintain the simplicity of use found in p-y curves, but they allow for the direct consideration of complex aspects of soil-structure interaction on the macroscopic response of the system, such as the progressive degradation of the soil’s stiffness and resistance, friction between the pile and the ground, and the hysteretic damping response of the soil, among others.
In this section, a macroelement capable of replicating the horizontal undrained cyclic response in terms of force-displacement of a monopile in a saturated clay layer is presented. The macroelement results from training a TANN on a numerical database obtained from finite element simulations of the macroscopic system.
@@
The TANN is combined with Proper Orthogonal Decomposition and applied to a macroscopic geotechnical system. Eventually, the studied 3D system is condensed into a macroelement with one degree of freedom.
@@ 4.3.2 Data generation and training of the TANN
The soil-pile system is reduced to a macroelement with one degree of freedom, considering the horizontal displacement of the point at the pile head positioned on the pile axis, $U_{H}$, as a function of the resultant of the horizontal forces at the pile head, $F_{H}$.
@@
As already shown, the threshold of 1e-4 allows to achieve satisfactory results both in terms of accuracy in the reproduced response of the 1D equivalent system and of the reconstruction error of the microscopic fields.
4.3.3 Inference of the TANN
For this kind of applications the possibility of reconstructing the fields used as internal coordinates is particularly appealing. One of the major drawbacks of the classical macroelement formulations is the impossibility of retrieving microscopic information, once the macroscopic behavior has been simulated. This poses clear difficulties in the engineering design, since often structural components are encapsulated by the macro-material whose response is of interest. Thus, an additional model is needed to link the global response of the classical macroelement to that of the structural components. As shown above, this in not needed with the proposed POD-TANN approach.
@@ 4.3.3 Inference of the TANN
Figure 13: TANN predictions in inference of a strain path outside the training set for cycles of positive and negative amplitude (a) and asymmetric cycles (b).
@@
4.3.4 Discussion on the POD-TANN macroelement calibration
The proposed application of the POD-TANN approach to macroelement derivation demonstrates the significant potential of the methodology. However, its current implementation exhibits limitations.
A primary limitation of the approach lies in its specificity. The entire procedure, encompassing the computational setup, data collection, preprocessing, and analysis, is highly tailored to the particular scenario under investigation. When the scenario changes - whether in terms of domain geometry, material properties, or boundary conditions - the entire computational process must be repeated. This also includes the task of training a new neural network, which is intrinsically linked to the specific problem it was trained on. Consequently, the approach is restricted in its ability to generalize across different scenarios of the same problems without substantial reconfiguration and retraining.
Traditionally, the calibration of macroelements in geomechanics has been rooted in the parametric formulation of these elements, aimed at reducing the need for extensive numerical computations by expressing macroelement behavior in terms of adjustable parameters functions of scenario-specific geometric data and material properties (see, e.g., [40]).
In this context, Transfer Learning (TL) presents a promising strategy to extend the applicability of the POD-TANN approach across various scenarios.
Specifically, TL could enhance the POD-TANN methodology in two significant ways:
  • 1.
    Fine-tuning pre-trained networks: TL allows for the adaptation of a pre-trained neural network to new scenarios, even when these scenarios differ in domain geometry or material parameters. The pre-trained network, which encapsulates the learned behavior from the original scenario, can be fine-tuned using a smaller, scenario-specific dataset. This approach significantly reduces the computational effort associated with training a new network from scratch for each new problem. This concept has been successfully demonstrated in recent studies, such as [41], where TL has been used to adapt models to different geotechnical conditions efficiently.
  • 2.
    Development of parametric macroelement networks: TL algorithms can be designed to link specific model input parameters, such as geometry and material properties, to each new scenario. By incorporating these parameters into the force vector and initial conditions, it is possible to develop a parametric macroelement network. Such a network would generalize across a range of scenarios, providing a computationally efficient tool for geomechanical modeling. In this perspective, [42] demonstrated the potential of TL in parametric modeling by applying it to the stability analysis of slopes with varying spatial distributions, highlighting the achievable efficiency gain.
An important consideration in this context is the number of POD modes used in the model, as this directly influences the dimensionality of the internal state variable vector. The selection of POD modes is crucial, as it determines how well the reduced-order model can capture the essential features of the system. For TL to be effective across different scenarios, the selected number of POD modes should be representative across a range of macro-model parameters, including variations in geometry, material properties, and boundary conditions, as stated before. This requirement would ensures that the fixed number of POD modes (i.e. of ISV vector components) remains unchanged and sufficient for new scenarios, thereby enabling the successful application of TL, without TANN’s hyperparameters changes (e.g., the number of inputs, and consequently the number and density of hidden layers).
In this view, the need to maintain a consistent number of POD modes across different scenarios could serve as an additional criterion for determining the appropriate number of POD modes. Rather than solely focusing on ensuring that the learning targets are met, this approach would also consider the need to encompass a class of possible macro-model parameters, facilitating more effective TL.
This dual consideration could help balance the trade-off between model accuracy and computational efficiency while broadening the applicability of the POD-TANN approach.